OpenStudy (anonymous):

Find the domain and the equation of the vertical asymptote for the function f(x) = log(3 - x)

OpenStudy (anonymous):

Vertical asymptote is when x is not equal to 0

OpenStudy (anonymous):

@Loser66 @Hero

hero (hero):

Have you tried graphing log(3 - x)?

OpenStudy (anonymous):

I did on my calculator. I set x - 3 > 0 for domain and x -3 = 0, the way it is written is log(3 - x) but I think it is easier to change it to log(x-3)

OpenStudy (anonymous):

Domain: x > 3 (but the answer is -oo, 3), and I don't see why -oo would be included in the domain V.A. x = 3

OpenStudy (anonymous):

@Hero, I was thinking that log is positive values so it wouldn't include negative infinity

hero (hero):

The domain of log( ) occurs when the argument of log( ) is greater than zero. In this case the argument of log( ) is 3 - x, therefore 3 - x > 0.

OpenStudy (anonymous):

Oh so I can't just change it to x - 3 because I thought 3 -x and x -3 were the same thing according to wolfram alpha when I wanted to see if they could be switched around

hero (hero):

\(3 - x \ne x - 3\) Example: let x = 2 \(3 - 2 = 2 - 3\) \(1 = -1\) False

hero (hero):

|3 - x| = |x - 3|

OpenStudy (anonymous):

That is because of absolute value

hero (hero):

That's the only way they become equal.

OpenStudy (anonymous):

So to find the domain, I am setting 3 -x > 0 and 3 -x = 0 to find the vertical asymptote.

hero (hero):

Yes, that appears to be correct.

OpenStudy (anonymous):

3 - x > 0 Subtract 3 from both sides -x > -3

hero (hero):

No, I wouldn't approach it that way. 3 - x > 0 Add x to both sides: 3 > x

OpenStudy (anonymous):

Oh I would have got into some trouble had I solved it that way... that is where I need help is in my approaches to problems

OpenStudy (anonymous):

Ok, so I get 3 > x for the domain and 3 = x for the equation of the asymptote

OpenStudy (anonymous):

Actually, x = 3 is the equation

hero (hero):

You can solve it your way as well. I just do it the way I do because I avoid negatives at all costs.

hero (hero):

Because if you have negatives you end up having to flip signs and in my mind, flipping a sign is not a valid mathematical step.

OpenStudy (anonymous):

That is a good rule to follow. The domain would be (-oo, 0) U (0, 3)

hero (hero):

The domain is (-oo, 3)

OpenStudy (anonymous):

Why is it -oo? Is that because a logarithm has a domain of -oo, oo?

OpenStudy (anonymous):

We said that the domain is 3 > x

hero (hero):

Yes and 3 > x is equivalent to x < 3

hero (hero):

Did you not learn this when you were little?

OpenStudy (anonymous):

Oh I was thinking of 3 > x as the same as x > 3

hero (hero):

Try not to get mixed up with that

OpenStudy (anonymous):

No, I get confused on this concept of switching signs

hero (hero):

Whether you have 3 > x or x < 3, the pointy end is always facing x and bigger end is always facing 3.

hero (hero):

That's how you keep up with things without confusing them.

OpenStudy (anonymous):

I'm sure I heard that when I was little but I forgot it.

OpenStudy (anonymous):

I'm going to put that in my notes.